Problem 14-26 (LO 14-2, 14-4, 14-6, 14-9)

In the early part of 2018, the partners of Hugh, Jacobs, and Thomas sought assistance from a local accountant. They had begun a new business in 2017 but had never used an accountant’s services.

Hugh and Jacobs began the partnership by contributing $170,000 and $120,000 in cash, respectively. Hugh was to work occasionally at the business, and Jacobs was to be employed full-time. They decided that year-end profits and losses should be assigned as follows:

  • Each partner was to be allocated 10 percent interest computed on the beginning capital balances for the period.
  • A compensation allowance of $5,000 was to go to Hugh with a $27,000 amount assigned to Jacobs.
  • Any remaining income would be split on a 4:6 basis to Hugh and Jacobs, respectively.

  

In 2017, revenues totaled $195,000, and expenses were $158,000 (not including the partners’ compensation allowance). Hugh withdrew cash of $10,000 during the year, and Jacobs took out $15,000. In addition, the business paid $6,500 for repairs made to Hugh’s home and charged it to repair expense.

On January 1, 2018, the partnership sold a 15 percent interest to Thomas for $58,000 cash. This money was contributed to the business with the bonus method used for accounting purposes.

1. What journal entries should the partnership have recorded on December 31, 2017?

A. Record entry to reclassify payment made to repair personal residence.

B. Record entry to close drawings accounts for 2017.

C. Record entry to close revenue and expense accounts for 2017.

D. Record the distribution of net income to partners.

2. What journal entry should the partnership have recorded on January 1, 2018?

A. Record the payment made by Thomas using the bonus method.
Answer #1

Please find below table useful to compute desired results: -

C D Dr. Amt. Cr. Amt. Working Notes: - Jacob Hugh act 6500 =G14*10/100 =H14*10/100 6500 Total 43500 =G4+H4 =13-14 =G6+H6 =15-